OpenStudy (jennychan12):

antiderivative of (x^2+4x)^1/3 ?

OpenStudy (jennychan12):

\[\sqrt[3]{x^2+4x}\]

OpenStudy (jennychan12):

i know that u = x^2+4x so du = (2x+4)dx where do i go from there?

OpenStudy (jennychan12):

@agent0smith @Dido525

OpenStudy (anonymous):

well u multipl that by 1/3 too

OpenStudy (anonymous):

1/3 dx (x^2 + 4x) = 1/3* (2x + 4)..I think so

OpenStudy (jennychan12):

how did you get that? can u explain?

OpenStudy (jennychan12):

cuz du = (2x+4)dx = (2(x+2))dx so 1/2du = (x+2)dx then...?

OpenStudy (anonymous):

actually, the value of power multplies..hmm du = 1/3*(2x + 4)dx => 3*du = (2x + 4)dx

zepdrix (zepdrix):

I'm pretty sure you need to either do a Trig-Sub, or Integration-By-Parts might also be possible. No U-Sub available though :( Have you learned either of those methods yet?

OpenStudy (jennychan12):

yeah the substitution method.

zepdrix (zepdrix):

Trig subs? :D

OpenStudy (jennychan12):

i think so? remind me what that is?

OpenStudy (kainui):

I think you might want to add and subtract +4 and -4 under the cube root so that you can factor.

OpenStudy (anonymous):

google the concept behind it :)

OpenStudy (beginnersmind):

idk how to do this, I'm just posting so that I get notified when someone answers it.

OpenStudy (kainui):

I could reduce it down so that it becomes \[2^{3/5}\int\limits_{}^{}\tan^{3/5} \theta \sec \theta d \theta\] after making the substitution: \[(x+2)=2\sec \theta\]

OpenStudy (kainui):

the fraction should be 5/3 not 3/5 on the 2 and tan(theta)

OpenStudy (anonymous):

this link will help you :) http://www.the-mathroom.ca/freebs/cali3/cali3.htm

OpenStudy (kainui):

I can't seem to crack it. You're in calculus 1 or 2?

OpenStudy (jennychan12):

1

OpenStudy (jennychan12):

where did the tangent, secant stiff come from?

zepdrix (zepdrix):

Ah sorry I was busy :C Kanoo been helping you? c:

OpenStudy (kainui):

I was doing trigonometric substitution, but you don't learn that until cal 2. If there's an easier way with just using u-sub then I would be surprised! Good luck zeppy! =P

zepdrix (zepdrix):

Ya I dunno sorry Jenny :c This doesn't seem like a problem you should be doing yet, hmm.

OpenStudy (kainui):

Check this out: http://www.wolframalpha.com/input/?i=(x%5E2%2B4x)%5E1%2F3dx&t=crmtb01 I've never even heard of that stuff. Hypergeometric function? lol

zepdrix (zepdrix):

Woah that looks fun :O

OpenStudy (jennychan12):

oy.....! -_-

OpenStudy (anonymous):